OpenStudy (solomonzelman):

Is the best approach here to apply the power rule to solve the bottom integral and then split the fraction and solve each of the 3 integrals separately? \[\int\limits_{~~ }^{~~} ~\frac{x^3-x^2+1}{\int\limits_{ }^{ } x^2}~dx\]

OpenStudy (solomonzelman):

\[\int\limits_{~~ }^{~~} ~\frac{x^3-x^2+1}{\int\limits_{ }^{ } x^2}~dx\]

OpenStudy (sidsiddhartha):

yeah thats a good way

OpenStudy (solomonzelman):

I am getting \[\int\limits\limits_{~~ }^{~~} ~\frac{3(x^3-x^2+1)}{ x^3}~dx\] then I apply the sum rule, (I'll draw it) ..

OpenStudy (anonymous):

i have never seen an integral over another integral

OpenStudy (solomonzelman):

yes, it is quite wierd, they just make it confusing....